OpenStudy (anonymous):

Math help: For the line 3x-y=2, the x intercept is ____ and they intercept is ____. A. -2;2/3 B. 2/3;-2 C. 2;-2/3 D. -2/3;2

OpenStudy (anonymous):

To find x -intercept put y = 0 in our equation 3x - y = 2 so we get 3x - 0 = 2 that is 3x = 2 , x = 2/3 now to find y - intercept put x = 0 in equation so we get 0 - y = 2 , y = -2

OpenStudy (anonymous):

Is it clear to you ?

OpenStudy (anonymous):

yea, thanks!

OpenStudy (anonymous):

You are welcome :)